OpenStudy (anonymous):

Find the slope of line passing through the points (4,8) and (6, -2). A. 1/5 B. 5 C. -1/5 D. -5

OpenStudy (anonymous):

@luckythebest

OpenStudy (anonymous):

D

OpenStudy (anonymous):

how did you get D ? @luckythebest is it 4-6/8-(-2)

OpenStudy (anonymous):

@ilovedaivd you put the change in x in numerator :P its change in y/change in x; not change in x/change in y

OpenStudy (anonymous):

\[\frac{ 2+8 }{ 4-6 }= -5\]

OpenStudy (anonymous):

so 4-(-2)/8-6 ???

OpenStudy (anonymous):

That is an easy way http://www.wolframalpha.com/input/?i=%284%2C8%29+and+%286%2C+-2%29+slope+

OpenStudy (anonymous):

\[(X _{1},Y _{1}),(X _{2},Y _{2})\]

OpenStudy (anonymous):

\[ m= \frac{ Y _{2} -Y _{1}}{ X _{2}-X _{1} }\]